# FIPER: Generalizable Factorized Fields for Joint Image Compression and Super-Resolution

### v2

@@ 3.2 Super-Resolution with Factorized Fields
We represent a super-resolved image using our Factorized Fields, where coefficients and basis are generated by networks $F_{\text{coeff}}$ and $F_{\text{basis}}$ from a low-resolution Image:
@@
$\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{i}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (4)
@@
where $c^{\text{LR}}_{i}(x)=\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))_{i}(x)$ and $b^{\text{LR}}_{i}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))_{i}(\gamma_{i}(x))$. Note that we sample the outputs $\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively.
@@
Our model comprises three main components: Coefficient Backbone, Basis Swin Transformer, and Factorized Fields Reconstruction. As shown in  Fig. 3, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$. The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations, and fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$. The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  Eq. 4, where $s$ is the scale factor. In our experiments, $\gamma_{i}(x)$ is a Sawtooth function: $\gamma_{i}(x)=x\mod k_{i}$, with $k_{i}\in R$. We optimize model parameters using an $L_{1}$ loss function. To demonstrate our method’s effectiveness, we use existing SR methods [14, 15, 75] as the Coefficient Backbone. For the Basis Swin Transformer, we employ Swin Transformer Blocks [51] with a series of downsampling operations. We use a dilation-like downsampling technique( Fig. 2c) to accommodate the sawtooth sampling pattern. The final basis is refined using additional upsampling and convolution layers.
3.3 Image Compression with Factorized Fields and Super-Resolution
@@
Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality. However, as discussed in  Sec. 1,  Sec. 2, and Sec. 3.1, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models [38, 42], or decompression modules [63, 22], while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through explicit modeling of frequency components and Factorized Fields formulation as in  Eq. 3. In addition, with our trained SR model described in  Sec. 3.2, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data. Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline. In the following, we respectively present the proposed joint framework for the single and multi-image compression tasks.
3.3.1 Single Image Compression
@@ 4.1 Image Super-Resolution
Experimental Setup.
We conduct extensive experiments to validate the effectiveness of our Factorized Fields representation for Super-Resolution tasks. Following the strategy outlined in [13] and [18], we adapt the same-task pretraining approach for all the Super-Resolution models. Unlike these previous works, we leverage the SA-1B dataset from [35], which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet [20], which is used for [13] and [18] pretraining. Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction. We further conduct experiments to compare training performance with SA-1B and ImageNet in  Tab. 1. While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
@@
For pretraining, we utilize AdamW optimizer with learning rate 1e-4, batch size 16, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning. Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input. As for the hyperparameter of Factorized Fields, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information. More implementation details can be found in supplementary materials.
Quantitative Results.
Tab. 1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods. We evaluate the methods using five benchmark datasets, including Set5 [5], Set14 [74], BSD100 [55], Urban100 [31], and Manga109 [56]. For quantitative metrics, PSNR and SSIM are reported. The average relative improvement of 204.4% in PSNR over the baseline across the five datasets is calculated using the formula $(c-b)/(a-b)$, where $b$ represents the PSNR of a representative baseline, SwinIR, $a$ represents the PSNR of SOTA, HAT-L, and $c$ represents the PSNR of our best model, HAT-L-F. This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD [75], DAT [15], and HAT [14]—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively. These models exhibit significant improvements when compared to their counterparts ATD-L, DAT-L, and HAT-L, which possess similar parameter counts. It is important to note that only [14] provides a large-scale model. To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages. To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in [12, 14]. The results demonstrate that its performance remains consistent with only minor perturbations.
@@
Lastly, to evaluate the effectiveness of our Factorized Fields, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq. 3. Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output. Although this approach results in reduced performance, which indicates the representation does act as an imperative role in modeling image information, the Basis Swin Transformer with Sawtooth downsampling still contributes to improved reconstruction, even without Factorized-Fields decoding, highlighting its effectiveness.
Visual Comparison.
We provide the visual comparison in Fig. 6. The images are randomly sampled from the DIV2K dataset. Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
Multi-frame SR.
Due to the space limit, we provide additional comparisons on multi-frame SR in the supplementary materials.
@@
Xie (MM 21’) -0.78 2.93 6.00 50.0
Cheng (CVPR 22’) 5.44 1.98 4.69 29.6
STF (CVPR 22’) -4.31 0.14 0.13 99.9
ELIC (CVPR 22’) -7.24 0.07 0.09 36.9
TCM (CVPR 23’) -11.74 0.16 0.15 76.7
@@ 4.2 Single- and Multi-Image Compressions
Experimental Setup.
We evaluate our Factorized Fields representation for image compression tasks, comparing it against state-of-the-art methods. Following our Super-Resolution setting in Experiment Setup from  Sec. 4.1, we use the same set of SA-1B for training. To emphasize our representation, we initialize compression and decompression modules in Fig. 5 from pre-trained [48] and the SR Module from those in Sec. 4.1, we then train the pipeline end to end on 256x256 patches for 200k iterations, with AdamW optimizer [52] with learning rate 1e-5, batch size 16, betas (0.9, 0.99) and other parameters set to PyTorch default.
@@
Rate-Distortion Performance Comparison.
We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including [49], [11], [82], [73], [17], [2], [42], [32], [57], [4], [61], and [30]. The classical image compression codec, VVC [67], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig. 7.
@@
Our analysis reveals several significant advantages of the FIPER framework in image compression tasks. The approach demonstrates substantial improvements in compression efficiency across various benchmark datasets, including Kodak, CLIC, and Tecnick, indicating its broad applicability. Notably, the multi-image compression strategy shows particularly promising results for larger image sets, suggesting scalability benefits. Furthermore, our method maintains competitive latency while significantly improving compression performance and balancing efficiency and quality.
@@
Table 4: Comparison of improvements of Factorized Fields. $\psi$ and $\alpha$ are the same in Eq. 4.
@@ 4.3 Ablation Studies
Influence of SR Priors in Image Compression.
We conduct experiments with various configurations to verify the proposed image compression pipeline’s effectiveness. As shown in Tab. 3, we present the quantitative performance of models trained with different values of $\lambda$ in Eq. 5. Specifically, TCM-HAT refers to substituting our SR Module with the original HAT [14] in the pipeline illustrated in Fig. 5b. TCM-HAT-F represents our complete pipeline, while TCM-HAT-F-Scratch denotes the same pipeline but with the SR Module initialized randomly. Our results demonstrate that integrating SR priors with image compression improves performance, and our proposed representation further enhances results. This highlights the robustness of our Factorized Fields in capturing fine details in image regression tasks.
Effectiveness of Factorized Fields Design.
We conduct experiments to verify our modification of Factorized Fields, modified from Eq. 1 to Eq. 3. The quantitative performance reported on single-image regression is shown in Tab. 4, where each result is measured after 256 iterations. Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics. Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.

### v3

@@ 3.2 Super-Resolution with Factorized Features
We represent a super-resolved image using our Factorized Features, where coefficients and basis are generated by networks $F_{\text{coeff}}$ and $F_{\text{basis}}$ from a low-resolution Image:
@@
$\hat{I}_{\text{SR}}(x)=\mathcal{P}\Bigl{(}{\textbf{Concat}_{i=1}^{N}}_{j=1}^{K}\Bigl{\{}c_{ij}^{\text{LR}}(x)\odot\psi(\alpha_{j}\cdot b_{i}^{\text{LR}}(\gamma_{i}(x)))\Bigl{\}}\Bigl{)},$ (8)
@@
where $c^{\text{LR}}(x)=\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))(x)$ and $b^{\text{LR}}(x)=F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))(\gamma(x))$. Note that we sample the outputs $\text{Conv}(F_{\textbf{coeff}}(I_{\text{LR}}))$ and $F_{\textbf{basis}}(F_{\textbf{coeff}}(I_{\text{LR}}))$ with coordinates $x$ and $\gamma(x)$, respectively.
@@
Our model comprises three main components: Coefficient Backbone $F_{\textbf{coeff}}$, Basis Swin Transformer $F_{\textbf{basis}}$, and Factorized Features Reconstruction. As shown in  Fig. 3, the process begins with $I_{LR}\in\mathbb{R}^{3\times H\times W}$. The Coefficient Backbone extracts features $X_{\text{coeff}}\in\mathbb{R}^{C_{c}\times H_{c}\times W_{c}}$, which are then used to generate coefficients $c$ through convolution and pixel shuffle operations. Then, $X_{\text{coeff}}$ is fed into the Basis Swin Transformer to produce a multi-scale basis $b=\{b_{1},...,b_{N}\},\ b_{i}\in\mathbb{R}^{C_{b_{i}}\times H_{b_{i}}\times W_{b_{i}}}$. The coefficients and basis are combined to reconstruct $I_{SR}\in\mathbb{R}^{3\times sH\times sW}$ using  Eq. 8, where $s$ is the scale factor.
We optimize model parameters using $L_{1}$ loss.
To demonstrate the effectiveness of our method, we use existing SR methods [16, 17, 91] as the Coefficient Backbone. For the Basis Swin Transformer, we employ simple Swin Transformer Blocks [59] with a series of special downsampling operations. Specifically, we use a dilation-like downsampling technique( Fig. 2.c) to accommodate the sawtooth sampling pattern. The final basis is refined using additional upsampling and convolution layers.
3.3 Image Compression with Factorized Features
@@
Image Compression, at its core, is to strike a balance between the amount of information contained in the latent bits and the final image quality. However, as discussed in  Sec. 1 and  Sec. 2, most recent works draw emphasis on how to retrieve implicit elements through designing different architectures, such as analysis transforms and entropy models [39, 44], or decompression modules [75, 24], while this paper aims at addressing the representation itself to better capture the structural correlations and thus achieve better image quality through Factorized Features formulation as in  Eq. 7
Priors from SR. In addition, with our trained SR model described in  Sec. 3.2, it intuitively serves as a strong prior for information recovery, i.e., it contains extensive knowledge of how to reconstruct missing details and enhance image quality by leveraging learned patterns from the training data. Thus, since Super-Resolution and Image Compression share the core principle of reconstructing and enhancing image details from low-quality sources, we can effectively integrate this prior into the compression pipeline.
@@ 4.1 Image Super-Resolution
Experimental Setup. We conduct extensive experiments to validate the effectiveness of our Factorized Features representation for Super-Resolution tasks. Following the strategy outlined in [15] and [20], we adapt the same-task pretraining approach for all the Super-Resolution models. Unlike these previous works, we leverage the SA-1B dataset from [37], which includes approximately 10 million images, where we randomly sample 4 million for training, which is much less that of ImageNet [22], which is used for [15] and [20] pretraining. Note that we use SA-1B just for its content-rich and high-resolution images since SA-1B is designed for training segmentation models, whereas ImageNet focuses primarily on single-class prediction. We further conduct experiments to compare training performance with SA-1B and ImageNet in  Tab. 1. While the performances are mostly equivalent, we witness faster convergence with SA-1B during training.
@@
For pretraining, we utilize AdamW optimizer with learning rate 2e-4, batch size 32, betas (0.9, 0.99), and other parameters set to PyTorch default, while we use learning rate 1e-5 during finetuning. Throughout training, the input is randomly cropped to $256\times 256$ and bicubicly resized to $64\times 64$ for Coefficient Backbone input. As for the hyperparameter of Factorized Features, the number $N$ of coefficient and basis is set to 6, and the scalars $\alpha_{j}$ are set to {1, 4, 16, 64} since we want to capture both base frequency $\alpha_{j}=1$ and high frequency $\alpha_{j}=64$ information. More implementation details can be found in supplementary materials.
Quantitative Results. Tab. 1 presents the quantitative comparison between our approach and state-of-the-art (SoTA) methods. We evaluate the methods using five benchmark datasets, including Set5 [6], Set14 [90], BSD100 [63], Urban100 [33], and Manga109 [64]. For quantitative metrics, PSNR and SSIM are reported. The average relative improvement of 204.4% in PSNR over the baseline across the five datasets is calculated using the formula $(c-b)/(a-b)$, where $b$ represents the PSNR of a representative baseline, SwinIR, $a$ represents the PSNR of SOTA, HAT-L, and $c$ represents the PSNR of our best model, HAT-L-F. This formula measures the relative performance gain of our model compared to the gap between HAT-L and SwinIR.
To validate the effectiveness of our framework, we employ three different SoTA SR models—ATD [91], DAT [17], and HAT [16]—as the Coefficient Backbone in our pipeline, which we denote as ATD-F, DAT-F, and HAT-F, respectively. These models exhibit significant improvements when compared to their counterparts, ATD-L, DAT-L, and HAT-L, which possess similar parameter counts. It is important to note that only [16] provides a large-scale model. To maintain fairness, we scale up ATD and DAT to match this size and train them under the same configuration as our model, including both pretraining and fine-tuning stages. To ensure an equitable comparison with other methods, we further train another model, HAT-F-ImageNet, using ImageNet as the pretraining dataset, following the protocols outlined in [14, 16]. The results demonstrate that its performance remains consistent with only minor perturbations.
@@
Lastly, to evaluate the effectiveness of our Factorized Features, we trained a model named HAT-F-Concat, which does not apply the formulation in Eq. 7. Instead, it concatenates the basis and coefficient directly and decodes the resulting features to produce the output. Although this approach results in reduced performance, which indicates the representation does act as an imperative role in modeling image information, the Basis Swin Transformer with Sawtooth downsampling still contributes to improved reconstruction, even without Factorized-Fields decoding, highlighting its effectiveness.
Visual Comparison. We provide the visual comparison in Fig. 7. The images are randomly sampled from the DIV2K dataset. Our method faithfully reconstructs the image details, whereas the other approaches suffer from over-smoothing or hallucinating details absent in the ground truth.
Multi-frame SR. Due to space limitations, we provide additional comparisons on multi-frame SR in the supplementary materials.
@@
Xie (MM’21) -0.78 2.93 6.00 50.0
Cheng (CVPR’22) 5.44 1.98 4.69 29.6
STF (CVPR’22) -4.31 0.14 0.13 99.9
ELIC (CVPR’22) -7.24 0.07 0.09 36.9
TCM (CVPR’23) -11.74 0.16 0.15 76.7
CCA (NeurIps’24) -20.97 0.11 0.09 64.9
@@ 4.2 Single- and Multi-Image Compressions
Experimental Setup. We evaluate our Factorized Features for image compression tasks against state-of-the-art methods. Following the Super-Resolution setup in Sec. 4.1, we train on the SA-1B dataset. To emphasize our representation, compression and decompression modules are initialized from pre-trained models [55], then fine-tuned end-to-end on 256×256 patches for 200k iterations using the AdamW optimizer [60] (learning rate=1e-5, batch size=16, betas=(0.9, 0.99)).
@@
Rate-Distortion Performance Comparison. We compare our model with State-of-the-Art learned end-to-end image compression algorithms, including [56], [13], [98], [87], [19], [3], [44], [34], [66], [5], [71], and [32]. The classical image compression codec, VVC [80], is also tested by using VTM12.1.The rate-distortion performance on various datasets, including Kodak, Tecnick’s old test set with resolution 1200$\times$1200, and CLIC Professional Validation, is shown in Fig. 8.
@@
Our experiments validate the effectiveness of FIPER in image compression, demonstrating substantial improvements across benchmarks (Kodak, CLIC, Tecnick). The multi-image compression strategy particularly excels with larger image sets, highlighting strong scalability. Additionally, our method achieves competitive latency, significantly improving compression efficiency while balancing quality and speed.
Table 4: Comparison of improvements of Factorized Features. $\psi$ and $\alpha$ are the same in Eq. 8.
@@ 4.3 Ablation Studies
Influence of SR Priors in Image Compression. We conduct experiments with various configurations to verify our image compression pipeline’s effectiveness. TCM-HAT refers to using the original HAT [16] instead of our SR Module in Fig. 6b. TCM-HAT-F represents our full pipeline, while TCM-HAT-F-Scratch initializes the SR Module randomly. Results demonstrate that integrating SR priors improves image compression performance, and our representation further enhances results, highlighting the robustness of Factorized Fields.
Effectiveness of Factorized Features Design. We conduct experiments to verify our multi-frequency modulation. The quantitative performance reported on single-image regression is shown in Tab. 4, where each result is measured after 256 iterations. Compared to baseline results, our refinements in modeling pixel-level frequency have significantly improved all performance metrics. Additionally, our results demonstrate that the modulation function $\psi$ and the scalar $\alpha$ are interdependent, each essential to the other’s function.